The Pawlson Company's year-end balance sheet is shown below. Its cost of common equity is 18%, its before-tax cost of debt is 11%, and its marginal tax rate is 40%. Assume that the firm's long-term debt sells at par value. The firm’s total debt, which is the sum of the company’s short-term debt and long-term debt, equals $1,154. The firm has 576 shares of common stock outstanding that sell for $4.00 per share.

Assets Liabilities And Equity
Cash $ 120 Accounts payable and accruals $ 10
Accounts receivable 240 Short-term debt 54
Inventories 360 Long-term debt 1,100
Plant and equipment, net 2,160 Common equity 1,716
Total assets $2,880 Total liabilities and equity $2,880

Calculate Pawlson's WACC using market-value weights. Round your answer to two decimal places. Do not round your intermediate calculations.
%

Please Help!

Thank you!
Answer #1

Solution :

As per the information given in the question            

Market value of the Common Equity = No. of shares outstanding * Market price per share

= 576 * $ 4.00 = $ 2,304

Market Value of the debt = Par value of Debt = $ 1,154

Thus Weight of Common Equity = [ 2304 / ( 2304 + 1154) ] = 2304 / 3458 = 0.666281

Thus Weight of Debt = [ 1154 / ( 2304 + 1154) ] = 1154 / 3458 = 0.333719

Cost of Common equity = 18 %   ; Before tax cost of debt = 11 %

Calculation of WACC :

The formula for calculating the weighted average cost of capital is =

WACC = [ Ke * We ] + [ ( Kd * ( 1- t ) ) * Wd ]

Ke = Cost of Common equity ; We = Weight of Common equity ; Kd = Before tax Cost of debt    ; t = Income tax rate ; Wd = Weight of debt

As per the information available in the question we have

Ke = 18 % = 0.18   ; We = 0.666281 ;   Kd = 11 % = 0.11 ; t = 40 % = 0.40   ; Wd = 0.333719

Applying the above values in the formula we have

= [ 0.18 * 0.666281 ] + [ (0.11 * ( 1 – 0.40 ) ) * 0.333719 ]

= [ 0.18 * 0.666281 ] + [ (0.11 * 0.60 * 0.333719 ]

= [ 0.119931 + 0.022025   ]

= 0.141956

= 14.1956 %   

= 14.20 % ( when rounded off to four decimal places)

Thus Pawlson’s WACC based on Market value weights is = 14.20 %